Charlies Furniture Store has been in business for several years. The firms owners have described the store as a high-price
Answer #1

A. Current sales = turnover*total assets

= 0.4*800,000 =$320,000

ROI = Margin*turnover

= 32%*0.4

= 12.8%

B.turnover required = 12.8%/20% = 0.64

Sales required = 0.64*800,000 =$512,000

C.Actual increase in sales required =$192,000

I.e. 60%

D.when the prices are not reduced, margin will only be reduced by the amount spent on advertising since margin = net operating income/sales and advertising cost will lead to a reduction in net operating income

Turnover might increase based on campaign as it will lead to higher sales

ROI will not be affected much, it might reduce a little because of cost of campaign